OpenStudy (anonymous):

Simplify: x squared minus 10x + 9 over x squared + 2x minus 3

jhonyy9 (jhonyy9):

\[(x ^{2}-10x +9)/(x ^{2} +2x -3) =((x-9)(x-1))/((x+3)(x-1))=(x-9)/(x+3)\]